The road from A to B is 10km uphill followed by 20km downhill. A motorcyclist averages 36km/h uphill and 90km/h downhill. Calculate the average speed from
1. A to B 2. B to A

Answer by Alan3354(69443)

The road from A to B is 10km uphill followed by 20km downhill. A motorcyclist averages 36km/h uphill and 90km/h downhill.
-------
Assuming the up and down speeds are the same for both directions (which seems contradictory):
---
Calculate the average speed from
1. A to B
Time uphill = 10/36 hours
Time downhill = 20/90 hours
Total time = 1/2 hour
Total distance = 30 km
Avg speed = 30/(1/2) = 60 km/hr
=======================
2. B to A
Time uphill = 20/36 hours
Time downhill = 10/90 hours
Total time = 2/3 hour
Total distance = 30 km
Avg speed = 30/(2/3) = 45 km/hr
---------------------------------------
As pointed out by the other tutor, this won't work in the real world, only in a math problem.